Consider the matrices

A = −2 7 1

3 4 1

8 1 5,



B =8 1 5

3 4 1

−2 7 1,



C = −2 7 1

3 4 1

2 −7 3


Find elementary matrices E1, E2 and E3 such that


(5.1) E1A = B,

(5.2) E1B = A,

(5.3) E2A = C,

(5.4) E3C = A.




1
Expert's answer
2021-05-24T13:47:51-0400

"\textbf{Solution:}"

Let "E" be the identity matrix:

"E=\left(
\begin{array}{cccc}
1 & 0 & \ldots & 0\\
0 & 1 & \ldots & 0\\
\vdots & \vdots & \ddots & \vdots\\
0 & 0 & \ldots & 1
\end{array}
\right)=\left(
\begin{array}{cccc}
\ldots \\
E_{i} \\
\ldots \\
 E_{j} \\
\ldots
\end{array}
\right)".

Let "P_{ij}" be the permutation matrix:

"P_{ij}=\left(
\begin{array}{cccc}
\ldots \\
E_{j} \\
\ldots \\
 E_{i} \\
\ldots
\end{array}
\right)".

Then the result of left multiplication of the matrix "P_{ij}" by the matrix

"A=\left(
\begin{array}{cccc}
\ldots \\
A_{i} \\
\ldots \\
 A_{j} \\
\ldots
\end{array}
\right)"

is the matrix obtained from the original matrix "A" by permuting its "i"-th and "j"-th rows:

"P_{ij}A=\left(
\begin{array}{cccc}
\ldots \\
A_{j} \\
\ldots \\
 A_{i} \\
\ldots
\end{array}
\right)".

For example,

"P_{23}A=\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 0 & 1 \\
 0 & 1 & 0
\end{pmatrix}\begin{pmatrix}
 a_{11} & a_{12} & a_{13} \\
 a_{21} & a_{22} & a_{23} \\
 a_{31} & a_{32} & a_{33}
\end{pmatrix}=\begin{pmatrix}
 a_{11} & a_{12} & a_{13} \\
 a_{31} & a_{32} & a_{33} \\
 a_{21} & a_{22} & a_{23}
\end{pmatrix}".

--------------------------------------------------------------------------------------------------------------------

5.1) "E_{1}A=B"

"E_{1}\begin{pmatrix}
 -2\ & 7\ & 1 \\
 3 & 4\ & 1 \\
 8 &1\ & 5
\end{pmatrix}=\begin{pmatrix}
 8 & 1\ & 5 \\
 3 & 4\ & 1 \\
 -2\ & 7\ & 1
\end{pmatrix}".


"P_{ij}=P_{13}\Rightarrow E_{1}=P_{13}=\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0
\end{pmatrix}".

"\textbf{Check:}"

"\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0
\end{pmatrix}\begin{pmatrix}
 -2\ & 7\ & 1 \\
 3 & 4\ & 1 \\
 8 &1\ & 5
\end{pmatrix}=\begin{pmatrix}
 8 & 1\ & 5 \\
 3 & 4\ & 1 \\
 -2\ & 7\ & 1
\end{pmatrix}".

--------------------------------------------------------------------------------------------------------------------

5.2) "E_{1}B=A"

"E_{1}\begin{pmatrix}
 8 & 1\ & 5 \\
 3 & 4\ & 1 \\
 -2\ & 7\ & 1
\end{pmatrix}=\begin{pmatrix}
 -2\ & 7\ & 1 \\
 3 & 4\ & 1 \\
 8 &1\ & 5
\end{pmatrix}".


"P_{ij}=P_{13}\Rightarrow E_{1}=P_{13}=\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0
\end{pmatrix}".

"\textbf{Check:}"

"\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0
\end{pmatrix}\begin{pmatrix}
 8 & 1\ & 5 \\
 3 & 4\ & 1 \\
 -2\ & 7\ & 1
\end{pmatrix}=\begin{pmatrix}
 -2\ & 7\ & 1 \\
 3 & 4\ & 1 \\
 8 &1\ & 5
\end{pmatrix}".

--------------------------------------------------------------------------------------------------------------------

Let "A=\\left(\n\\begin{array}{cccc}\n\\ldots \\\\\nA_{i} \\\\\n\\ldots \\\\\n A_{j} \\\\\n\\ldots\n\\end{array}\n\\right), \t\n\\,\\widetilde{A}=\\left(\n\\begin{array}{cccc}\n\\ldots \\\\\nA_{i}+\\lambda A_{j} \\\\\n\\ldots \\\\\n A_{j} \\\\\n\\ldots\n\\end{array}\n\\right)", then


"\widetilde{A}=N_{ij}(\lambda)\left(
\begin{array}{cccc}
\ldots \\
A_{i} \\
\ldots \\
 A_{j} \\
\ldots
\end{array}
\right)=\left(
\begin{array}{cccc}
1 && 0 & \ldots & 0 & \ldots & 0 &\ldots & 0\\
0 && 1 & \ldots & 0 & \ldots & 0 &\ldots & 0\\
\vdots && \vdots & \ddots & \vdots & \ddots & \vdots & \ddots & \vdots\\
0 && 0 & \ldots & 1 & \ldots & \lambda & \ldots & 0\\
\vdots && \vdots & \ddots & \vdots & \ddots & \vdots & \ddots & \vdots\\
0 && 0 & \ldots & 0 & \ldots & 0 & \ldots & 1
\end{array}
\right)\left(
\begin{array}{cccc}
\ldots \\
A_{i} \\
\ldots \\
 A_{j} \\
\ldots
\end{array}
\right)=\left(
\begin{array}{cccc}
\ldots \\
A_{i}+\lambda A_{j} \\
\ldots \\
 A_{j} \\
\ldots
\end{array}
\right)." --------------------------------------------------------------------------------------------------------------------

5.3) "E_{2}A=C"

"E_{2}\begin{pmatrix}
 -2\ & 7\ & 1 \\
 3 & 4\ & 1 \\
 8 &1\ & 5
\end{pmatrix}=\begin{pmatrix}
 -2\ & 7 & 1 \\
 3 & 4 & 1 \\
 2 &-7\ & 3
\end{pmatrix}".

1-st occasion:

"\begin{pmatrix}
 2 & -7 & 3 
\end{pmatrix}=\begin{pmatrix}
 8 & 1 & 5 
\end{pmatrix}+\lambda\begin{pmatrix}
 -2 & 7 & 1 
\end{pmatrix}".

"\begin{cases}
 2=8-2\lambda, 
 \\
 -7=1+7\lambda,
 \\
 3=5+\lambda.
 \end{cases}"

"\lambda\in\varnothing".

2-nd occasion:

"\begin{pmatrix}
 2 & -7 & 3 
\end{pmatrix}=\begin{pmatrix}
 8 & 1 & 5 
\end{pmatrix}+\lambda\begin{pmatrix}
 3 & 4 & 1 
\end{pmatrix}".

"\begin{cases}
 2=8+3\lambda, 
 \\
 -7=1+4\lambda,
 \\
 3=5+\lambda.
 \end{cases}"

"\lambda=-2".

"i=3, j=2, \lambda=-2\Rightarrow N_{ij}(\lambda)=N_{32}(-2)."

"E_{2}=N_{32}(-2)=\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & -2 & 1
\end{pmatrix}".

"\textbf{Check:}"

"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & -2 & 1
\end{pmatrix}\begin{pmatrix}
 -2\ & 7\ & 1 \\
 3 & 4\ & 1 \\
 8 &1\ & 5
\end{pmatrix}=\begin{pmatrix}
 -2\ & 7 & 1 \\
 3 & 4 & 1 \\
 2 &-7\ & 3
\end{pmatrix}".

--------------------------------------------------------------------------------------------------------------------

5.4) "E_{3}C=A"

"E_{3}\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1\\
 2 & -7 & 3
\end{pmatrix}=\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 8 & 1 & 5
\end{pmatrix}".

1-st occasion:

"\begin{pmatrix}
 8 & 1 & 5 
\end{pmatrix}=\begin{pmatrix}
 2 & -7 & 3 
\end{pmatrix}+\lambda\begin{pmatrix}
 -2 & 7 & 1 
\end{pmatrix}".

"\begin{cases}
 8=2-2\lambda, 
 \\
 1=-7+7\lambda,
 \\
 5=3+\lambda.
 \end{cases}"

"\lambda\in\varnothing".

2-nd occasion:

"\begin{pmatrix}
 8 & 1 & 5 
\end{pmatrix}=\begin{pmatrix}
 2 & -7 & 3 
\end{pmatrix}+\lambda\begin{pmatrix}
 3 & 4 & 1 
\end{pmatrix}".

"\begin{cases}
 8=2+3\lambda, 
 \\
 1=-7+4\lambda,
 \\
 5=3+\lambda.
 \end{cases}"

"\lambda=2".

"i=3, j=2, \lambda=2\Rightarrow N_{ij}(\lambda)=N_{32}(2)."

"E_{3}=N_{32}(2)=\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 2 & 1
\end{pmatrix}".

"\textbf{Check:}"

"\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 2 & 1
\end{pmatrix}\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1\\
 2 & -7 & 3
\end{pmatrix}=\begin{pmatrix}
 -2 & 7 & 1 \\
 3 & 4 & 1 \\
 8 & 1 & 5
\end{pmatrix}".

-----------------------------------------------------------------------------------------------------------------

"\textbf{Answer:}"

"\boxed{E_{1}=\begin{pmatrix}
 0 & 0 & 1 \\
 0 & 1 & 0 \\
 1 & 0 & 0
\end{pmatrix}, \, E_{2}=\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & -2 & 1
\end{pmatrix}, \, E_{3}=\begin{pmatrix}
 1 & 0 & 0 \\
 0 & 1 & 0 \\
 0 & 2 & 1
\end{pmatrix}}"